# Categorical Flow Matching on Statistical Manifolds

### v2

@@ 2.1 Information Geometry
[…]
@@ 2.2 Conditional Flow Matching on Riemannian Manifold
Consider a smooth Riemannian manifold $\mathcal{M}$ with the Riemannian metric $g$, a probability path $p_{t}:[0,1]\to\mathcal{P(M)}$ is a curve of probability densities over $\mathcal{M}$. A flow $\psi_{t}:[0,1]\times\mathcal{M}\to\mathcal{M}$ is a time-dependent diffeomorphism defined by a time-dependent vector field $u_{t}:[0,1]\times\mathcal{M}\to T\mathcal{M}$ via the ordinary differential equation (ODE): $\frac{d}{dt}\psi_{t}(x)=u_{t}(\psi_{t}(x))$. The flow matching objective directly regresses the vector field $u_{t}$ with a time-dependent neural net $v(x_{t},t)$ where $x_{t}:=\psi_{t}(x)$. However, this objective is generally intractable. Both [37, 14] demonstrated that a tractable objective can be derived by conditioning on the target data $x_{1}$ at $t=1$ of the probability path. The Riemannian flow matching objective can be formulated as [14]
@@
where $q$ is the data distribution, $p_{0}$ is the prior distribution, and $x_{t}:=\psi_{t}(x|x_{0},x_{1})$ denotes the conditional flow. [14] further demonstrated that if the exponential map and logarithm map can be evaluated in closed-form, the condition flow can be defined as $x_{t}=\exp_{x_{0}}(t\log_{x_{0}}x_{1})$, and the corresponding vector field can be calculated as $u_{t}(x_{t}|x_{0},x_{1})=\frac{d}{dt}x_{t}=\log_{x_{t}}(x_{1})/(1-t)$. We also adapt this formulation for our statistical flow, which we will elaborate on in the next section. We note that, since we are working with manifolds of probability measures $\mathcal{M}=\mathcal{P(X)}$, we will use $p,q$ for probability densities over the manifold $\mathcal{P(X)}$ and $\mu,\nu$ for probability measures (or probability masses for discrete distributions) on the manifold $\mathcal{P(X)}$ to avoid confusion.
@@ 3.1 Statistical Manifold of Categorical Distributions
$d_{\text{cat}}(\mu,\nu)=2\arccos\left(\sum_{i=1}^{n}\sqrt{\mu_{i}\nu_{i}}\right)$ (3)
@@
$\langle u,v\rangle_{\mu}=\sum_{i=1}^{n}\frac{u_{i}v_{i}}{\mu_{i}},\quad\mu\in\mathcal{P}^{\circ},u,v\in T_{\mu}(\mathcal{P})$ (4)
where $\mathcal{P}^{\circ}$ denotes the interior of the statistical manifold $\mathcal{P}$. Note that the inner product is ill-defined on the boundary, causing numerical issues near the boundary. To circumvent this issue, we introduce the following diffeomorphism
$\pi:\mathcal{P}\to S^{n-1}_{+},\quad\mu_{i}\mapsto x_{i}=\sqrt{\mu_{i}}$ (5)
@@
$d_{S}(\pi(\mu),\pi(\nu))=\frac{1}{2}d_{\mathrm{cat}}(\mu,\nu),\quad\mu,\nu\in\mathcal{P}$ (6)
A proof is provided in Appendix A.3). This indicates that we can work with the geodesic distance on the unit sphere instead with up to a constant factor:
$d_{S}(x,y)=\arccos(\langle x,y\rangle),\quad x,y\in S^{n-1}_{+}$ (7)
@@ 3.2 Statistical Flow Matching
$\mathcal{L}_{\text{SFM}}=\mathbb{E}_{t\sim U[0,1],x_{0}\sim\pi_{*}(p_{0}(\mu)),x_{1}\sim\pi_{*}(q(\mu))}\left[\|v(x_{t},t)-u^{S}_{t}(x_{t}|x_{0},x_{1})\|_{2}^{2}\right]$ (8)
@@
$\mathcal{L}_{\text{LinearFM}}=\mathbb{E}_{t\sim U[0,1],\mu_{0}\sim p_{0}(\mu),\mu_{1}\sim q(\mu)}\left[\|v(\mu_{t},t)-(\mu_{1}-\mu_{0})\|_{2}^{2}\right]$ (9)
@@ 3.3 Optimization View of Statistical Flow Matching
We further provide an interpretation of our proposed statistical flow matching framework from the perspective of an optimization problem. From the optimization viewpoint, for a generative model on a statistical manifold, we want to minimize the discrepancy between the target distributions and the generated distributions. Naturally, the Kullback-Leibler divergence $D_{\text{KL}}(\mu(\theta)\|\mu(\theta_{1}))$ can be used as a measure of statistical distance. We note that the Fisher information metric can be obtained as the Hessian of the KL divergence with respect to the parameter $\theta$. This can be demonstrated by the fact that the KL divergence reaches the global minimum of 0 at $\theta=\theta_{1}$, so all first-order partial derivatives are zero and the Hessian is positive semi-definite. Therefore, Taylor expansion of KL divergence at $\theta_{1}$ with $\Delta\theta=\theta-\theta_{1}$ gives
@@
$=\frac{1}{2}\sum_{jk}\Delta\theta_{j}\Delta\theta_{k}g_{jk}(\theta_{1})=\frac{1}{2}\|\Delta\theta\|_{g}^{2}$
From the geometric viewpoint, the geodesic, by definition, is a (locally) length-minimizing curve with respect to the corresponding Riemannian metric. Therefore, by following the direction of the vector field that decreases the geodesic element $ds^{2}=\|d\theta\|^{2}_{g}\approx\|\Delta\theta\|^{2}_{g}$, we are indeed following the steepest direction that minimizes the local KL divergence. In this sense, the Fisher information metric defines a second-order optimization scheme for the KL divergence by following the “steepest” direction of the Hessian. Indeed, The steepest direction $\Delta\theta$ that decreases the KL divergence is known as the natural gradient in the existing literature [2, 3] and has been explored in optimization known as natural gradient descent[47, 41, 45]. Instead of optimizing along the normal gradient, the natural gradient is defined as ~$\tilde{\nabla}\mathcal{L}=F^{-1}\nabla\mathcal{L}$ where $F$ is the Fisher information matrix estimated from a batch of sampled data. Results established in [2, 41] demonstrated that stochastic natural gradient descent is asymptotically “Fisher efficient” and is indeed the steepest descent direction in the manifold of distributions where distance is measured in small local neighborhoods by the KL divergence. Following the geodesic defined by the Fisher information metric, our SFM framework also shares these benefits with an additional advantage of analytical expressions for geodesics, as we focus on the family of categorical distributions instead of general statistical models. Such a theoretical connection may contribute to the better performance of SFM.
@@ 3.5 Exact Likelihood Calculation
$\frac{\partial}{\partial t}\log p_{t}(x_{t})+\mathrm{div}_{g}(v_{t})(x_{t})=0$ (11)
@@
[…]
where the trajectory $x_{t}$ is obtained by solving the differential equation $\frac{\partial}{\partial t}x_{t}=v_{t}(x_{t})$ reverse through time with the initial condition $x_{1}$ at $t=1$. In this way, we have $\log p(x_{1})=\log p^{\text{ODE}}+\log p_{0}(x_{0})$ where $\log p_{0}(x_{0})$ is the log-likehood of the prior distribution at data point $x_{0}$. In practice, we follow previous work [7] to use Hutchinson’s trace estimator [29] to efficiently obtain an unbiased estimation of the divergence using standard normal random variables. To further account for the transform $\pi$ from $\mathcal{P}$ to $S^{n-1}_{+}$ and the reverse transform $\pi^{-1}$, additional change of measure formulae need to be applied. Consider the pushforward of the probability measure $P$ over $\mathcal{P}$ under the diffeomorphism $\pi$ defined by $\pi_{*}P(V):=P(\pi^{-1}V)$, we have the change of measure identity $\pi_{*}P(\pi(\mu))|\det d\pi(\mu)|=P(x)$ for $x=\pi(\mu)$. Therefore, by adding the two log-determinant of the pushforward measures, the log-likelihood can be formulated as
$\log p_{1}(\mu_{1})=\log|\det d\pi^{-1}(x_{1})|+\log p^{\text{ODE}}(x_{1})+\log|\det d\pi(\mu_{0})|+\log p_{0}(\mu_{0})$ (13)
@@
$\log p(\delta)\geq\mathbb{E}_{q(\tilde{\mu}_{1}|\delta)}[-\log q(\tilde{\mu}_{1}|\delta)+\log p(\delta|\tilde{\mu}_{1})+\log p_{1}(\tilde{\mu}_{1})]$ (14)
@@ 4.2 Binarized MNIST
We extend our SFM to model discrete generation tasks in computer vision. The binarized MNIST dataset [52] is the binarized version of the original MNIST dataset [33] by thresholding the original continuous value to be either 0 or 1, thus can be viewed as a 2-class generation task with a data dimension of $78428^{2}=$ 784. We also compared D3PM [6] with an additional absorbing state. We used a convolutional neural net adopted from [56] with additional additive Fourier-based time embeddings. The quantitative evaluation of NLL and Fréchet inception distance (FID) are shown in Tab.1.
Table 1: NLL and FID of different discrete models on binarized MNIST. * is from [7].
SFM w/ OT SFM w/o OT LinearFM D3PM DDSM DirichletFM
@@ 5 Related Work
As we put a special interest in discrete generation, we start with the existing work on discrete generation. We first list a few traditional autoregressive models [16, 49, 63] and masked language models [18, 53] but will skip them as we mainly focus on the diffusion and flow matching models. Existing diffusion and flow-based discrete generative models can be categorized into two main groups. The first group relies on stochastic jumps of Markov chains by treating discrete classes as Markov states. By choosing a proper transition matrix, the target one-hot distribution can be gradually noised into the stationary distribution. Noticeably, this approach can also adopt an additional absorbing state to mimic the mask token in masked language modeling [6]. D3PM [6] adapted the discrete-time Markov chain with the diffusion framework and SEDD [38] proposed a more efficient training scheme by learning the score entropy between different states. [11, 54] extended it to the continuous-time Markov chain by using the infinitesimal generator (rate matrix) instead, and [12] further applied the flow matching framework. Although the transition matrix or the rate matrix determines the entire diffusion trajectory, there is no canonical way of choosing it for different generation tasks to control the mixing rate. Also, due to the presence of discrete jumps of the Markov chain, exact likelihood calculation is no longer feasible. Only variational bounds were derived in [6, 12]. The other group directly works with the probability simplex or the logit space with certain assumptions of the underlying geometric structure [35]. As an example, our linear flow matching assumes a Euclidean geometry on the probability simplex and is often used as a baseline in previous work [59]. The multinomial diffusion [27] assumed a Euclidean geometry on the logits space so interpolation became multiplication back on the probability simplex. Dirichlet diffusion (DDSM) [7] and Dirichlet flow matching (DirichletFM) [59] exerted a strong prior assumption on the probability path as the Jacobi diffusion process. We provide a more detailed analysis of these models in Appendix A.4. For models that directly operate on the logit space, the Bayesian flow network (BFN) assumed Gaussian distributions on the logit space with Bayesian update rules. [39] also assumed a Euclidean geometry in the logit space with targets as soft one-hot logits. The assumptions made in these models may not always hold, e.g., for the Swiss roll on simplex dataset in Fig.3. These assumptions also did not necessarily capture the true geometry of the underlying statistical manifold. In contrast, our proposed SFM framework does not exert any strong prior assumptions and is aware of the intrinsic geometry by following the geodesics.

### v3

@@ 2.1 Information Geometry
In this work, we are interested in learning a parameterized family of probability measures. It is known from information theory that all probability measures over the sample space form the structure known as statistical manifold. Mathematically, consider probability densities $p=\frac{\mathop{}\!\mathrm{d}\mu}{\mathop{}\!\mathrm{d}\nu}:\mathcal{X}\to\mathbb{R}$ defined by the Radon-Nikodym derivative where $\mu$ is a probability measure on the sample space $\mathcal{X}$ and $\nu$ is the reference measure on $\mathcal{X}$. Suppose the statistical manifold $\mathcal{P}=\mathcal{P}(\mathcal{X})=\{p:\int_{\mathcal{X}}\mathop{}\!\mathrm{d}\mu=\int_{\mathcal{X}}p(x;\theta)\mathop{}\!\mathrm{d}\nu=1\}$ is parameterized by $\theta=(\theta_{1},\theta_{2},\dots,\theta_{n})\in\Theta$, this parameterization naturally provides a coordinate system for $\mathcal{P}$ on which each point is a probability measure $\mu$ with the corresponding probability density function $p(x;\theta)$. The Fisher information metric is defined as
$g_{jk}(\theta)=\mathbb{E}_{X}\left[\frac{\partial\log p(X;\theta)}{\partial\theta_{j}}\frac{\partial\log p(X;\theta)}{\partial\theta_{k}}\right]=\int_{\mathcal{X}}\frac{\partial\log p(x;\theta)}{\partial\theta_{j}}\frac{\partial\log p(x;\theta)}{\partial\theta_{k}}p(x;\theta)\mathop{}\!\mathrm{d}\nu.$ (1)
Rao demonstrated in his seminal paper [51] that statistical manifold can be equipped with the Fisher information metric to obtain a Riemannian structure, the study of which is known as information geometry [5, 4, 8]. This geometric view of statistical manifolds allows us to derive key geometric concepts for our statistical flow matching framework. For example, a geodesic $\gamma(t):[0,1]\to\mathcal{P}$ defines a “shortest” path (under the Riemannian metric) connecting two probability measures on the statistical manifold. The geodesic distance between two probability measures, also known as the Fisher-Rao distance [51], measures the similarity between them. The tangent space $T_{\mu}(\mathcal{P})$ at a point $\mu\in\mathcal{P}$ can be naturally identified with the affine subspace $T_{\mu}(\mathcal{P})=\{\nu|\int_{\mathcal{X}}\mathop{}\!\mathrm{d}\nu=0\}$ where each element $\nu$ is a signed measure over $\mathcal{X}$. The exponential map $\exp_{\mu}:T_{\mu}(\mathcal{P})\to\mathcal{P}$ and logarithm map $\log_{\mu}:\mathcal{P}\to T_{\mu}(\mathcal{P})$ can also be defined on the statistical manifold. While the geodesic for a parameterized family of probability measures can be obtained numerically by solving the geodesic equation when closed-form expression is unknown (see Appendix A.1), it usually requires expensive simulations. Fortunately, closed-form geodesic distances are available for many common distributions including categorical, multinomial, and normal distributions [44], which motivates our method.
@@ 2.2 Conditional Flow Matching on Riemannian Manifold
Consider a smooth Riemannian manifold $\mathcal{M}$ with the Riemannian metric $g$, a probability path $p_{t}:[0,1]\to\mathcal{P(M)}$ is a curve of probability densities over $\mathcal{M}$. A flow $\psi_{t}:[0,1]\times\mathcal{M}\to\mathcal{M}$ is a time-dependent diffeomorphism defined by a time-dependent vector field $u_{t}:[0,1]\times\mathcal{M}\to T\mathcal{M}$ via the ordinary differential equation (ODE): $\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\psi_{t}(x)=u_{t}(\psi_{t}(x))$. The flow matching objective directly regresses the vector field $u_{t}$ with a time-dependent neural net $v(x_{t},t)$ where $x_{t}:=\psi_{t}(x)$. However, this objective is generally intractable. Both [38, 14] demonstrated that a tractable objective can be derived by conditioning on the target data $x_{1}$ at $t=1$ of the probability path. The Riemannian flow matching objective can be formulated as [14]
@@
where $q$ is the data distribution, $p_{0}$ is the prior distribution, and $x_{t}:=\psi_{t}(x|x_{0},x_{1})$ denotes the conditional flow. [14] further demonstrated that if the exponential map and logarithm map can be evaluated in closed-form, the condition flow can be defined as $x_{t}=\exp_{x_{0}}(t\log_{x_{0}}x_{1})$, and the corresponding vector field can be calculated as $u_{t}(x_{t}|x_{0},x_{1})=\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}x_{t}=\log_{x_{t}}(x_{1})/(1-t)$. We also adapt this formulation for our statistical flow, which we will elaborate on in the next section. We note that, since we are working with manifolds of probability measures $\mathcal{M}=\mathcal{P(X)}$, we will use $p,q$ for probability densities over the manifold $\mathcal{P(X)}$ and $\mu,\nu$ for probability measures (or probability masses for discrete distributions) on the manifold $\mathcal{P(X)}$ to avoid confusion.
@@ 3.1 Statistical Manifold of Categorical Distributions
$d_{\text{cat}}(\mu,\nu)=2\arccos\left(\sum_{i=1}^{n}\sqrt{\mu_{i}\nu_{i}}\right).$ (3)
@@
$\langle u,v\rangle_{\mu}=\sum_{i=1}^{n}\frac{u_{i}v_{i}}{\mu_{i}},\quad\mu\in\mathcal{P}_{+},u,v\in T_{\mu}(\mathcal{P})$ (4)
where $\mathcal{P}_{+}$ denotes the statistical manifold of positive categorical distributions. Note that the inner product is ill-defined on the boundary, causing numerical issues near the boundary. To circumvent this issue, we introduce the following diffeomorphism
$\pi:\mathcal{P}\to S^{n-1}_{+},\quad\mu_{i}\mapsto x_{i}=\sqrt{\mu_{i}},$ (5)
@@
$d_{S}(\pi(\mu),\pi(\nu))=\frac{1}{2}d_{\mathrm{cat}}(\mu,\nu),\quad\mu,\nu\in\mathcal{P}.$ (6)
A proof is provided in Appendix A.3. This indicates that we can work with the geodesic distance on the unit sphere instead with up to a constant factor:
$d_{S}(x,y)=\arccos(\langle x,y\rangle),\quad x,y\in S^{n-1}_{+}.$ (7)
@@ 3.2 Statistical Flow Matching
$\mathcal{L}_{\text{SFM}}=\mathbb{E}_{t\sim U[0,1],x_{0}\sim\pi_{*}(p_{0}(\mu)),x_{1}\sim\pi_{*}(q(\mu))}\left[\|v(x_{t},t)-u^{S}_{t}(x_{t}|x_{0},x_{1})\|_{2}^{2}\right],$ (8)
@@
$\mathcal{L}_{\text{LinearFM}}=\mathbb{E}_{t\sim U[0,1],\mu_{0}\sim p_{0}(\mu),\mu_{1}\sim q(\mu)}\left[\|v(\mu_{t},t)-(\mu_{1}-\mu_{0})\|_{2}^{2}\right].$ (9)
@@ 3.3 Optimization View of Statistical Flow Matching
We further provide an interpretation of our proposed statistical flow matching framework from the perspective of an optimization problem. From the optimization viewpoint, for a generative model on a statistical manifold, we want to minimize the discrepancy between the target distributions and the generated distributions. Naturally, the Kullback-Leibler divergence $D_{\text{KL}}(\mu(\theta)\|\mu(\theta_{1}))$ can be used as a measure of statistical distance. We note that the Fisher information metric can be obtained as the Hessian of the KL divergence with respect to the parameter $\theta$. This can be demonstrated by the fact that the KL divergence reaches the global minimum of 0 at $\theta=\theta_{1}$, so all first-order partial derivatives are zero, and the Hessian is positive semi-definite. Therefore, Taylor expansion of KL divergence at $\theta_{1}$ with $\Delta\theta=\theta-\theta_{1}$ gives
@@
$=\frac{1}{2}\sum_{jk}\Delta\theta_{j}\Delta\theta_{k}g_{jk}(\theta_{1})=\frac{1}{2}\|\Delta\theta\|_{g}^{2}.$
From the geometric viewpoint, the geodesic, by definition, is a (locally) length-minimizing curve with respect to the corresponding Riemannian metric. Therefore, by following the direction of the vector field that decreases the geodesic element $\mathop{}\!\mathrm{d}s^{2}=\|\mathop{}\!\mathrm{d}\theta\|^{2}_{g}\approx\|\Delta\theta\|^{2}_{g}$, we are indeed following the steepest direction that minimizes the local KL divergence. In this sense, the Fisher information metric defines a second-order optimization scheme for the KL divergence by following the “steepest” direction of the Hessian. Indeed, The steepest direction $\Delta\theta$ that decreases the KL divergence is known as the natural gradient in the existing literature [2, 3] and has been explored in optimization known as natural gradient descent[48, 42, 46]. Instead of optimizing along the normal gradient, the natural gradient is defined as ~$\tilde{\nabla}\mathcal{L}=F^{-1}\nabla\mathcal{L}$ where $F$ is the Fisher information matrix estimated from a batch of sampled data. Results established in [2, 42] demonstrated that stochastic natural gradient descent is asymptotically “Fisher efficient” and is indeed the steepest descent direction in the manifold of distributions where distance is measured in small local neighborhoods by the KL divergence. Following the geodesic defined by the Fisher information metric, our SFM framework also shares these benefits with an additional advantage of analytical expressions for geodesics, as we focus on the family of categorical distributions instead of general statistical models. Such a theoretical connection may contribute to the better performance of SFM.
@@ 3.5 Exact Likelihood Calculation
$\frac{\partial}{\partial t}\log p_{t}(x_{t})+\mathrm{div}_{g}(v_{t})(x_{t})=0,$ (11)
@@
$\log p^{\text{ODE}}(x_{1})=\int_{1}^{0}\mathrm{div}_{g}(v_{t})(x_{t})\mathop{}\!\mathrm{d}t,$ (12)
where the trajectory $x_{t}$ is obtained by solving the differential equation $\frac{\partial}{\partial t}x_{t}=v_{t}(x_{t})$ reverse through time with the initial condition $x_{1}$ at $t=1$. In this way, we have $\log p(x_{1})=\log p^{\text{ODE}}+\log p_{0}(x_{0})$ where $\log p_{0}(x_{0})$ is the log-likehood of the prior distribution at data point $x_{0}$. In practice, we follow previous work [7] to use Hutchinson’s trace estimator [30] to efficiently obtain an unbiased estimation of the divergence using standard normal random variables. To further account for the transform $\pi$ from $\mathcal{P}$ to $S^{n-1}_{+}$ and the reverse transform $\pi^{-1}$, additional change of measure formulae need to be applied. Consider the pushforward of the probability measure $P$ over $\mathcal{P}$ under the diffeomorphism $\pi$ defined by $\pi_{*}P(V):=P(\pi^{-1}V)$, we have the change of measure identity $\pi_{*}P(\pi(\mu))|\det\mathop{}\!\mathrm{d}\pi(\mu)|=P(x)$ for $x=\pi(\mu)$. Therefore, by adding the two log-determinant of the pushforward measures, the log-likelihood can be formulated as
$\log p_{1}(\mu_{1})=\log|\det\mathop{}\!\mathrm{d}\pi^{-1}(x_{1})|+\log p^{\text{ODE}}(x_{1})+\log|\det\mathop{}\!\mathrm{d}\pi(\mu_{0})|+\log p_{0}(\mu_{0}).$ (13)
@@
$\log p(\delta)\geq\mathbb{E}_{q(\tilde{\mu}_{1}|\delta)}[-\log q(\tilde{\mu}_{1}|\delta)+\log p(\delta|\tilde{\mu}_{1})+\log p_{1}(\tilde{\mu}_{1})],$ (14)
@@ 4.2 Binarized MNIST
We extend our SFM to model discrete generation tasks in computer vision. The binarized MNIST dataset [53] is the binarized version of the original MNIST dataset [34] by thresholding the original continuous value to be either 0 or 1, thus can be viewed as a 2-class generation task with a data dimension of $78428^{2}=$ 784. We also compared D3PM [6] and DFM [22] as examples of masked diffusion models for discrete generation. We used a convolutional neural net adopted from [57] with additional additive Fourier-based time embeddings. The quantitative evaluation of NLL and Fréchet inception distance (FID) are shown in Tab.1.
Table 1: NLL and FID of different discrete models on binarized MNIST. The NLLs in the parenthesis are discrete NLLs; therefore, they are not directly comparable. * is from [7].
Model SFM w/ OT SFM w/o OT SFM w/o $\pi$ LinearFM
[…]
Model DirichletFM DDSM D3PM DFM
@@ 5 Related Work
As we put a special interest in discrete generation, we start with the existing work on discrete generation. We first list a few traditional autoregressive models [16, 50, 64] and masked language models [18, 54] but will skip them as we mainly focus on the diffusion and flow matching models. Existing diffusion and flow-based discrete generative models can be categorized into two main groups. The first group relies on stochastic jumps of Markov chains by treating discrete classes as Markov states. By choosing a proper transition matrix, the target one-hot distribution can be gradually noised into the stationary distribution. Noticeably, this approach can also adopt an additional absorbing state to mimic the mask token in masked language modeling [6]. D3PM [6] adapted the discrete-time Markov chain with the diffusion framework and SEDD [39] proposed a more efficient training scheme by learning the score entropy between different states. [11, 55] extended it to the continuous-time Markov chain by using the infinitesimal generator (rate matrix) instead, and [12, 22] further applied the flow matching framework. Although the transition matrix or the rate matrix determines the entire diffusion trajectory, there is no canonical way of choosing it for different generation tasks to control the mixing rate. Also, due to the presence of discrete jumps of the Markov chain, exact likelihood calculation is no longer feasible. Only variational bounds were derived in [6, 12]. The other group directly works with the probability simplex or the logit space with certain assumptions of the underlying geometric structure [36]. As an example, our linear flow matching assumes a Euclidean geometry on the probability simplex and is often used as a baseline in previous work [60]. The multinomial diffusion [28] assumed a Euclidean geometry on the logits space so interpolation became multiplication back on the probability simplex. Dirichlet diffusion (DDSM) [7] and Dirichlet flow matching (DirichletFM) [60] exerted a strong prior assumption on the probability path as the Jacobi diffusion process. We provide a more detailed analysis of these models in Appendix A.4. For models that directly operate on the logit space, the Bayesian flow network (BFN) assumed Gaussian distributions on the logit space with Bayesian update rules. [40] also assumed a Euclidean geometry in the logit space with targets as soft one-hot logits. The assumptions made in these models may not always hold, e.g., for the Swiss roll on simplex dataset in Fig.3. These assumptions also did not necessarily capture the true geometry of the underlying statistical manifold. In contrast, our proposed SFM framework does not exert any strong prior assumptions and is aware of the intrinsic geometry by following the geodesics.